# Robust parameter estimation within minutes on gravitational wave signals from binary neutron star inspirals

### v1

@@ III.1 Accuracy of ripple waveforms
We verify the correctness of our waveform generators by comparing them against the LALsuite LIGO Scientific Collaboration et al. (2018) implementation. This is done by computing the mismatch $M$ between the waveforms generated by the two implementations, which is defined by
@@
In Fig. 2, the distribution of the mismatch between the ripple waveforms against the corresponding LALsuite implementation is shown. Specifically, we have sampled 1000010000{10000}10000 parameters from the distributions shown in Tab. 1. The mismatch is computed on a frequency grid in the range $[20,2048]$ Hz with spacing $\Delta f=1/T$, where we take the duration $T$ to be $128$ s, and with the design sensitivity of Advanced LIGO Aasi et al. (2015) used for the PSD. For comparison, we also show the mismatches for ripple’s BBH waveform IMRPhenomD for the same parameter ranges, excluding the tidal deformabilities. All ripple waveform samples have mismatches $M\lesssim 10^{-8}$ against their LALsuite counterpart, making them indistinguishable for parameter estimation Pürrer and Haster (2020).
@@
Component masses $[0.5M_{\odot},3M_{\odot}]$
Component aligned spins $[-0.05,0.05]$
@@ III.2 Injection-recovery test
To demonstrate the robustness of our PE pipeline, we perform injection-recovery tests and report the results in a percentile-percentile (p-p) plot Cook et al. (2006). We inject simulated GW signals into realizations of noise from a detector network and conduct PE on the simulated data with Jim. After performing a batch of such analyses, we calculated the credible level at which each true parameter appears in its marginal posterior distribution. If the pipeline delivers unbiased estimates of the source properties, it is expected that the true parameters occur in the $x\%$ credible interval $x\%$ of the time. Therefore, the cumulative distribution of the credible level should trend along the diagonal, which is often checked visually with a p-p plot.
@@
$\mathcal{M}$ detector-frame chirp mass $[M_{\odot}]$ $[0.88,2.61]$ $[1.18,1.21]$ $[1.485,1.490]$
@@
$\chi_{i}$ aligned spins $[-0.05,0.05]$ $[-0.05,0.05]$ $[-0.05,0.05]$
$\Lambda_{i}$ dimensionless tidal deformabilities $[0,5000]$ $[0,5000]$ $[0,5000]$
$d_{L}$ luminosity distance $[\rm{Mpc}]$ $[30,300]$ $[1,75]$ $[1,500]$
$t_{c}$ coalescence time $[\rm{s}]$ $[-0.1,0.1]$ $[-0.1,0.1]$ $[-0.1,0.1]$
$\phi_{c}$ coalescence phase $[0,2\pi]$ $[0,2\pi]$ $[0,2\pi]$
$\cos\iota$ cosine of inclination angle $[-1,1]$ $[-1,1]$ $[-1,1]$
$\psi$ polarization angle $[0,\pi]$ $[0,\pi]$ $[0,\pi]$
$\alpha$ right ascension $[0,2\pi]$ $[0,2\pi]$ $[0,2\pi]$
$\sin\delta$ sine of declination $[-1,1]$ $[-1,1]$ $[-1,1]$
@@
For both waveforms considered here, we create 100 GW events by sampling source parameters uniformly over the ranges given in Tab. 2. We reject sampled injection parameters that result in a signal-to-noise ratio below $12$. The duration is set to $128$ s, with a frequency range of $[20,2048]$ Hz. The synthetic GW signal is injected into a network consisting of the Advanced LIGO Aasi et al. (2015) and Advanced Virgo Acernese et al. (2015) detectors at their design sensitivities. We evolve 10001000{1000}1000 Markov chains and stop the flowMC training phase once the NF achieves an average acceptance rate of $20\%$. Instead of the individual aligned spins, we show the effective spin $\chi_{\rm eff}$, defined by
@@
The resulting p-p plots are shown in Fig. 3. Since the plots generally trend along the diagonal, we conclude that our pipeline is robust.
Refer to captionRefer to caption
Figure 3: p-p plot for the injections done with TaylorF2 (left) and IMRPhenomD_NRTidalv2 (right), each created from 100 injections. For both of the waveform models, the p-p plots agree well with the diagonal, demonstrating the robustness of the Jim pipeline on analyzing BNS signals. The shaded regions indicate the $1\sigma$, $2\sigma$ and $3\sigma$ confidence levels due to finite sample sizes.
@@ IV Analysis on GW170817 and GW190425
With the robustness of our pipeline demonstrated, we apply Jim to the two BNS events observed so far: GW170817 Abbott et al. (2017) and GW190425 Abbott et al. (2020) [^3] While GW170817 has been confidently verified to be a BNS merger, there have been proposals that GW190425 could be a neutron-star–black-hole merger Foley et al. (2020); Han et al. (2020); Kyutoku et al. (2020); Hinderer et al. (2019); Coughlin and Dietrich (2019).. For both events, we take the publicly available data from GWOSC Abbott et al. (2021, 2023) as input. The priors used in the analysis are reported in Tab. 2.
For comparison, we have conducted the PE with both TaylorF2 and IMRPhenomD_NRTidalv2 using both Jim and Parallel-Bilby (pBilbySmith et al. (2020) to verify the accuracy of Jim. As mentioned in Sec. II.3, Jim utilizes relative binning, employing evosax to search for the reference parameters. We 10001000{1000}1000 Markov chains and stop the flowMC training phase once the NF achieves an average acceptance rate of $10\%$. The resulting posterior samples are compared in Sec.IV.1, and the run times are compared in Sec.IV.2.
@@ IV.1 Posterior comparison
In Figures 4, 5, 6 and 7 in Appendix B, we show the visual comparison between the posterior samples obtained from Jim and pBilby and find agreement between the posteriors. To make a quantitative statement, we report the Jensen-Shannon (JS) divergences Lin (1991) in bits between the marginal distributions in Tab. 3. Since the JS divergences lie between 00 and $1$ and the highest JS divergence is $0.01847$ bits ($0.014136$ bits) for GW170817 (GW190425), we conclude that the results of the Jim and pBilby pipelines are consistent with each other.
@@
Parameter TF2 NRTv2 TF2 NRTv2
$\mathcal{M}$ $0.001725$ $0.000516$ $0.003557$ $0.001884$
$q$ $0.005212$ $0.007894$ $0.004837$ $0.003883$
$\chi_{1}$ $0.005633$ $0.004301$ $0.002794$ $0.001904$
$\chi_{2}$ $0.003030$ $0.002671$ $0.002416$ $0.001400$
$\Lambda_{1}$ $0.001062$ $0.002208$ $0.008556$ $0.000652$
$\Lambda_{2}$ $0.000559$ $0.002186$ $0.005808$ $0.001450$
$d_{L}$ $0.001544$ $\bm{0.01847}$ $0.001273$ $0.009157$
$\phi_{c}$ $0.003500$ $0.010714$ $0.003338$ $0.007874$
$\cos\iota$ $0.001615$ $0.012851$ $0.006400$ $0.009502$
$\psi$ $0.004048$ $0.011036$ $0.001516$ $0.005820$
$\alpha$ $\bm{0.014008}$ $0.001258$ $\bm{0.009822}$ $\bm{0.014136}$
$\sin\delta$ $0.009570$ $0.001761$ $0.008934$ $0.006362$
Table 3: Jensen-Shannon divergences (in bits) between the marginal posterior obtained for GW170817 and GW190425 using TaylorF2 and IMRPhenomD_NRTidalv2 with Jim and pBilby, with the highest value of each comparison in bold. The JS divergences are bound between $[0,1]$.
@@ IV.2 Wall time
For the analysis of the real events, the wall time of Jim additionally includes the time spent on calculating the reference parameters with evosax to initialize the relative binning likelihood, which takes between $5$ and $10$ minutes. Most of the wall time is spent on the training of the NF, which takes around $15$ minutes for the real events, after which the final production samples are produced in less than a minute. We remark that the sampling time for GW170817 with the IMRPhenomD_NRTidalv2 waveform is slightly higher, since $10$ times more bins were used in the relative binning method to achieve a reliable approximation, due to the breakdown of relative binning as discussed in Sec. II.3.1.
@@
For the injected events, we report the median value for the runtime of the PE runs performed with both waveforms. Since we set the reference parameters to the injected parameters, the runtime no longer includes any runtime from running the evosax algorithm. We note that the wall time for simulated events is slightly higher than that of real events. This is because we have adjusted the sampling settings to be more robust to accommodate the broadly distributed simulated events in terms of signal-to-noise ratio and chirp masses. While the evaluation of the IMRPhenomD_NRTidalv2 waveform is slower than the TaylorF2 waveform, we note that the speed of Jim is mainly determined by the convergence rate of the NF, the value for the stopping criterion chosen by the user and the number of bins used in relative binning.
@@
We note that Ref. Wong et al. (2023a) achieved wall times of a few minutes for analyzing BBH mergers with Jim. In our case, the wall time is higher, partly due to the increased dimensionality of the problem by including the tidal deformabilities, but mainly because of different sampler settings. In particular, our settings aim for robustness rather than speed by training the NF until it achieves an average acceptance rate of $10\%$ ($20\%$) for the real events (injections). On the other hand, the settings chosen in Ref. Wong et al. (2023a) resulted in an average acceptance rate of $2\%$$5\%$ for the NF.
@@
Finally, for the real events, the wall times are compared against their equivalent pBilby, Bilby with relative binning (RB-Bilby Krishna et al. (2023)) [^4] The reference parameters are identical to the ones used by Jim., and Bilby with ROQ (ROQ-Bilby Morisaki et al. (2023)) [^5] Due to the unavailability of the ROQ bases of TaylorF2 with tidal effects and IMRPhenomD_NRTidalv2, the bases of IMRPhenomPv2_NRTidalv2 is used., in Tab. 4. ROQ is a technique that creates an efficient representation of GW approximants with a reduced set of basis elements that still accurately reconstruct the entire model space, thereby reducing the number of terms to be computed. These runs make use of nested sampling Skilling (2006) with $1024$, $1000$ and $1000$ live points for pBilby, RB-Bilby and ROQ-Bilby, respectively. The pBilby runs are intended to verify the accuracy of our pipeline. However, pBilby is known to be expensive; therefore, we use $10$ Intel Skylake Xeon Platinum $8174$ CPUs ($480$ cores in total) per run so that the PE completes in a reasonable amount of time. For RB-Bilby and ROQ-Bilby, we instead use a single Intel Xeon Silver $4310$ Processor CPU, such that $24$ cores are used per run. [^6] There are $12$ physical cores on an Intel Xeon Silver $4310$ Processor CPU. On the machines used, there are two such CPUs installed. When running with 24 cores requested, the CPUs dynamically switches between with and without hyper-threading Intel Corporation . For a conservative estimate of the energy consumption, we take only the thermal design power (TDP) of one CPU in Sec. V.3.
@@
($1$ GPU) ($480$ cores) ($24$ cores) ($24$ cores)
[…]
NRTv2 $(5.69+28.02)$ min $10.99$ h $4.68$ h $1.65$ h
[…]

### v2

@@ III.1 Accuracy of ripple waveforms
We verify the correctness of our waveform generators by comparing them against the LALsuite lalsuite implementation. This is done by computing the mismatch $M$ between the waveforms generated by the two implementations, which is defined by
@@
In Fig. 2, the distribution of the mismatch between the ripple waveforms against the corresponding LALsuite implementation is shown. Specifically, we have sampled 10000{10000} parameters from the distributions shown in Tab. 1. The mismatch is computed on a frequency grid in the range $[20,2048]$ Hz with spacing $\Delta f=1/T$, where we take the duration $T$ to be $128$ s, and with the design sensitivity of Advanced LIGO LIGOScientific:2014pky used for the PSD. The mismatches for the IMRPhenomD_NRTidalv2 waveform are computed with the Planck taper still included in the ripple implementation. For comparison, we also show the mismatches for ripple’s BBH waveform IMRPhenomD for the same parameter ranges, excluding the tidal deformabilities. All ripple waveform samples have mismatches $M\lesssim 10^{-8}$ against their LALsuite counterpart, making them indistinguishable for parameter estimation Purrer:2019jcp.
@@
Component masses $0.5M_{\odot},3M_{\odot}]$
Component aligned spins [0.05,0.05][-0.05,0.05]
@@ III.2 Injection-recovery test
To demonstrate the robustness of our PE pipeline, we perform injection-recovery tests and report the results in a percentile-percentile (p-p) plot cook2006validation. We inject simulated GW signals into realizations of noise from a detector network and conduct PE on the simulated data with Jim. After performing a batch of such analyses, we calculated the credible level at which each true parameter appears in its marginal posterior distribution. If the pipeline delivers unbiased estimates of the source properties, it is expected that the true parameters occur in the $x\%$ credible interval $x\%$ of the time. Therefore, the cumulative distribution of the credible level should trend along the diagonal, which is often checked visually with a p-p plot.
@@
$\mathcal{M}$ detector-frame chirp mass $M_{\odot}] [$0.88,2.61][0.88,2.61] $[1.18,1.21]$ $[1.485,1.490]$
@@
$\chi_{1},\chi_{2}$ aligned spins [0.05,0.05][-0.05,0.05] [0.05,0.05][-0.05,0.05] [0.05,0.05][-0.05,0.05]
$\Lambda_{1},\Lambda_{2}$ dimensionless tidal deformabilities $[0,5000]$ $[0,5000]$ $[0,5000]$
$d_{L}$ luminosity distance $\rm{Mpc}] [$30,300][30,300] $[1,75]$ $[1,500]$
$t_{c}$ coalescence time $\rm{s}] [$0.1,0.1][-0.1,0.1] [0.1,0.1][-0.1,0.1] [0.1,0.1][-0.1,0.1]
$\phi_{c}$ coalescence phase $0,2\pi] [$$0,2\pi] [$$0,2\pi]$
$\cos\iota$ cosine of inclination angle [1,1][-1,1] [1,1][-1,1] [1,1][-1,1]
$\psi$ polarization angle $0,\pi] [$$0,\pi] [$$0,\pi]$
$\alpha$ right ascension $0,2\pi] [$$0,2\pi] [$$0,2\pi]$
$\sin\delta$ sine of declination [1,1][-1,1] [1,1][-1,1] [1,1][-1,1]
@@
For both waveforms considered here, we create 100 GW events by sampling source parameters uniformly over the ranges given in Tab. 2. We reject sampled injection parameters that result in a signal-to-noise ratio below 1212.333In principle, this introduces a small bias due to a difference between the prior distributions used for injection and recovery. Since we have found this to be a minor effect, we neglect this bias in this work. The duration is set to $128$ s and the frequency range to $[20,2048]$ Hz. We use $100$ bins for relative binning. The synthetic GW signal is injected into a network consisting of the Advanced LIGO LIGOScientific:2014pky and Advanced Virgo VIRGO:2014yos detectors at their design sensitivities. We evolve 1000{1000} Markov chains and stop the flowMC training phase once the NF achieves an average acceptance rate of $20\%$. Instead of the individual aligned spins, we show the effective spin $\chi_{\rm eff}$, defined by
@@
The resulting p-p plots are shown in Fig. 3. The combined $p$-values are $0.74$ and $0.85$, for the injections made with the TaylorF2 and IMRPhenomD_NRTidalv2 waveforms, respectively. Since the p-p plots also trend along the diagonal, we conclude that our pipeline is robust.
Figure 3: p-p plot for the injections done with TaylorF2 (left) and IMRPhenomD_NRTidalv2 (right), each created from 100 injections. For both of the waveform models, the p-p plots agree well with the diagonal, demonstrating the robustness of the Jim pipeline on analyzing BNS signals. The shaded regions indicate the $1\sigma$, $2\sigma$, and $3\sigma$ credible levels due to finite sample size.
@@ IV Analysis on GW170817 and GW190425
With the robustness of our pipeline demonstrated, we apply Jim to the two BNS events observed so far: GW170817 LIGOScientific:2017vwq and GW190425 LIGOScientific:2020aai [^4] While GW170817 has been confidently verified to be a BNS merger, there have been proposals that GW190425 could be a neutron-star–black-hole merger Foley:2020kus; Han:2020qmn; Kyutoku:2020xka; Hinderer:2018pei; Coughlin:2019kqf.. For both events, we take the publicly available data from GWOSC LIGOScientific:2019lzm; KAGRA:2023pio as input. The priors used in the analysis are reported in Tab. 2.
For comparison, we have conducted the PE with both TaylorF2 and IMRPhenomD_NRTidalv2 using both Jim and Parallel-Bilby (pBilbySmith:2019ucc to verify the accuracy of Jim. As mentioned in Sec. II.3, Jim utilizes relative binning, employing evosax to search for the reference parameters. We evolve 1000{1000} Markov chains and stop the flowMC training phase once the NF achieves an average acceptance rate of $10\%$. We use $200$ bins ($1000$ bins), when using the TaylorF2 (IMRPhenomD_NRTidalv2) waveform, respectively. The resulting posterior samples are compared in Sec. IV.1, and the run times are compared in Sec. IV.2.
@@ IV.1 Posterior comparison
In Figures 4, 5, 6 and 7 in Appendix B, we show the visual comparison between the posterior samples obtained from Jim and pBilby and find agreement between the posteriors. To make a quantitative statement, we report the Jensen-Shannon (JS) divergences lin1991divergence in bits between the marginal distributions in Tab. 3. The JS divergences lie between 0 and $1$ and the highest JS divergence is $0.026363$ bits ($0.010634$ bits) for GW170817 (GW190425). To get a sense of these numbers, we can compare them against JS divergences of two numerical experiments. First, in the final column of Tab. 3, we show the maximal JS divergence for each parameter found after varying the seed of the random number generator (RNG) for the analysis of GW190425 with TaylorF2 $10$ times. This provides a lower bound on the expected JS divergence when comparing posterior distributions, as we expect that fluctuations in the RNG will only marginally affect the posteriors. Second, Ref. Romero-Shaw:2020owr proposed the criterion to consider two empirical distributions to be identical if their JS divergence is below $0.0022$ bits. While some of the JS divergences in Tab. 3 are above this threshold, we can conclude that there is agreement between the results of the Jim and pBilby pipelines.
@@
TF2 NRTv2 TF2 NRTv2 RNG
$\mathcal{M}$ $0.001725$ $0.001553$ $0.003557$ $0.001626$ $0.000261$
$q$ $0.005212$ $0.003418$ $0.004837$ $0.003102$ $0.000242$
$\chi_{1}$ $0.005633$ $0.003815$ $0.002794$ $0.001512$ $0.000491$
$\chi_{2}$ $0.003030$ $0.003373$ $0.002416$ $0.001533$ $0.000247$
$\Lambda_{1}$ $0.001062$ $0.002802$ $0.008556$ $0.000690$ $0.000363$
$\Lambda_{2}$ $0.000559$ $0.003725$ $0.005808$ $0.001395$ $0.000477$
$d_{L}$ $0.001544$ $\bm{0.026363}$ $0.001273$ $0.008755$ $0.000969$
$\phi_{c}$ $0.003500$ $0.013961$ $0.003338$ $0.007819$ $0.000281$
$\cos\iota$ $0.001615$ $0.022125$ $0.006400$ $0.009580$ $0.000289$
$\psi$ $0.004048$ $0.015805$ $0.001516$ $0.005477$ $0.000203$
$\alpha$ $\bm{0.014008}$ $0.001896$ $\bm{0.009822}$ $\bm{0.010634}$ $0.001624$
$\sin\delta$ $0.009570$ $0.000735$ $0.008934$ $0.008896$ $0.000640$
Table 3: Jensen-Shannon divergences (in bits) between the marginal posterior obtained for GW170817 and GW190425 using TaylorF2 and IMRPhenomD_NRTidalv2 with Jim and pBilby, with the highest value of each comparison in bold. The JS divergences are bound between $[0,1]$. The final column shows the maximal deviation one obtains from fluctuations in the random number generator (RNG).
@@ IV.2 Wall time
For the analysis of the real events, the wall time of Jim additionally includes the time spent on calculating the reference parameters with evosax to initialize the relative binning likelihood, which takes between $2.5$ and $6$ minutes. Most of the wall time is spent on the training of the NF after which the final production samples are produced in less than a minute.
@@
As a comparison, we also provide the runtimes of the injection-recovery tests of Sec. III in Tab. 4. For each set of injections, we report the median value of the runtimes. Since we set the reference parameters to the injected parameters, the runtime no longer includes the time spent on the evosax algorithm. We note that the median wall time for the simulated events analyzed with TaylorF2 is slightly higher than that of the real events. This is because we have adjusted the sampling settings to be more robust to accommodate the broadly distributed simulated events in terms of signal-to-noise ratio and chirp masses. While the evaluation of the IMRPhenomD_NRTidalv2 waveform is slower than the TaylorF2 waveform, we note that the speed of Jim is mainly determined by the convergence rate of the NF, the value for the stopping criterion chosen by the user and the number of bins used in relative binning.
@@
We note that Ref. Wong:2023lgb achieved wall times of a few minutes for analyzing BBH mergers with Jim. In our case, the wall time is higher, partly due to the increased dimensionality of the problem by including the tidal deformabilities, but mainly because of different sampler settings. In particular, our settings aim for robustness rather than speed by training the NF until it achieves an average acceptance rate of $10\%$ ($20\%$) for the real events (injections). On the other hand, the settings chosen in Ref. Wong:2023lgb resulted in an average acceptance rate of $2\%$$5\%$ for the NF.
@@
For the real events, the wall times are compared against their equivalent pBilby, Bilby with relative binning (RB-Bilby Krishna:2023bug), and Bilby with ROQ (ROQ-Bilby Morisaki:2023kuq) [^5] Due to the unavailability of the ROQ bases of TaylorF2 with tidal effects and IMRPhenomD_NRTidalv2, the bases of IMRPhenomPv2_NRTidalv2 are used., in Tab. 4. ROQ is a technique that creates an efficient representation of GW approximants with a reduced set of basis elements that still accurately reconstruct the entire model space, thereby reducing the number of terms to be computed. These runs make use of nested sampling Skilling:2006gxv, in particular using dynesty Speagle:2019ivv, employing $1024$, $1000$ and $1000$ live points for pBilby, RB-Bilby and ROQ-Bilby, respectively. For the RB-Bilby runs, we use the same number of bins as the corresponding Jim runs, i.e., $200$ ($1000$) bins for the TaylorF2 (IMRPhenomD_NRTidalv2) waveform. This corresponds to setting the tunable $\epsilon$ parameter, as defined in Eq. (10) of Ref. Zackay:2018qdy, to $\epsilon=0.15$ ($\epsilon=0.03$). The reference parameters are set to those found by the evosax algorithm from the corresponding Jim runs and are therefore not computed independently by bilby.
The pBilby runs are intended to verify the accuracy of our pipeline. However, pBilby is known to be expensive; therefore, we use $10$ Intel Skylake Xeon Platinum $8174$ CPUs ($480$ cores in total) per run so that the PE completes in a reasonable amount of time. For RB-Bilby and ROQ-Bilby, we instead use a single Intel Xeon Silver $4310$ Processor CPU, such that $24$ cores are used per run. [^6] There are $12$ physical cores on an Intel Xeon Silver $4310$ Processor CPU. On the machines used, there are two such CPUs installed. When running with 24 cores requested, the CPUs dynamically switches between with and without hyper-threading intel_cpu_datasheet_4310. For a conservative estimate of the energy consumption, we take only the thermal design power (TDP) of one CPU in Sec. V.3.
@@
(11 GPU) ($480$ cores) ($24$ cores) ($24$ cores)
GW170817 TF2 $(4.85+15.33)$ min $\phantom{0}9.64$ h $3.8$ h
NRTv2 $(5.38+25.59)$ min $10.99$ h $4.11$ h $1.65$ h
GW190425 TF2 $(2.63+18.30)$ min $\phantom{0}8.18$ h $2.81$ h
NRTv2 $(3.26+21.20)$ min $\phantom{0}4.91$ h $2.42$ h $0.97$ h
Injection TF2 $\phantom{(0.000+}24.76\phantom{)}$ min
NRTv2 $\phantom{(0.000+}18.02\phantom{)}$ min
@@
IV.3 Sampler efficiency
Finally, we compare the efficiency of the different samplers by investigating the effective sample size (ESS) produced by the different pipelines averaged over the GW170817 and GW190425 runs. Since bilby makes use of nested sampling, the nested samples are rejection sampled to obtain posterior samples. Therefore, we take the total number of final posterior samples as ESS values for the bilby runs. On the other hand, Jim relies on MCMC sampling such that it directly produces posterior samples, which can however be correlated. For the Jim posteriors, we therefore compute the ESS using Geyer’s initial monotone sequence criterion geyer_SSE_ref1; geyer_SSE_ref2 as implemented in the Arviz package arviz_2019. We estimate the ESS for each parameter dimension and report the average across all parameters sampled during inference. The Jim posterior samples used for this calculation are taken only from flowMC’s production loop (i.e., after the NF has been trained and its weights are frozen) and we discard the training samples for this analysis. Since the production loop is computationally efficient, this phase took less than $1$ minute for the GW events considered here and can therefore be easily extended to produce more samples. The results are reported in Tab. 5. We note that Jim can achieve a larger ESS than other pipelines which can be attributed to the use of the global sampler. Indeed, since the NF proposes samples independent of the history of each Markov chain, each accepted sample coming from the NF is an effective sample. Therefore, the ESS depends strongly on the global acceptance rate achieved in the GW events considered here.
Effective sample size
Jim $4.1\times 10^{4}$
pBilby $6.5\times 10^{3}$
RB-Bilby $5.8\times 10^{3}$
ROQ-Bilby $7.4\times 10^{3}$
Table 5: Comparison of the effective sample size (ESS), averaged over the GW events considered in this work, between the different pipelines.